# Balancing Act: Prioritization Strategies for LLM-Designed Restless Bandit Rewards

### v4

@@ 1 Introduction
Unfortunately, handcrafting reward functions is often a challenging and time-consuming task for humans because of the complex relationship between reward functions and policy outcomes [12, 8, 16]. Further, the multiagent nature of the RMAB problem adds a new twist to the problem of reward design in RL: It becomes fundamentally multi-objective. Consider the above example prompt asking for the prioritization of two subpopulations. As these subpopulations may contain different agents, selecting a reward function will most likely involve trading off the interests of the low-income vs. older beneficiaries, making this a multi-objective problem. If this multi-objective nature is ignored, a selected reward function might heavily favor one of the two groups (e.g., leading to the allocation of many resources to low-income beneficiaries, and no resources to older ones). Another tradeoff that emerges because of the multiagent nature of the problem is between aligning with the preference prompt and impacting agents not referenced in the prompt: Changing the reward functions to align with a prompt can lead to drastic changes in the number of resources distributed to subpopulations not referenced in the user prompt (e.g., in the example prompt, the selected reward might lead to taking away all resources from well-educated beneficiaries). If not adequately balanced, this can cause strong side effects unintended by the user.
@@
To our knowledge, we are the first to address the multi-objective nature of LLM-powered reward design in RMABs in particular and multiagent planners in general. Closest to our paper is the work by Behari et al. [7] who proposed a fully LLM-based Decision-Language Model for RMABs to generate and select reward functions (as code) from human language prompts. However, as argued in Sections 2 and 4, the DLM model is not properly equipped to handle the multi-objective nature of the problem, as the LLM selects functions in an unpredictable, hard-to-control and sometimes (clearly) suboptimal way that does not adequately take into account and balance the different objectives.
@@
We present a Social Choice Language Model (SCLM) that designs reward functions (as Python code) aligned with complex, multi-objective human language preferences; see Figure 3 for an overview of SCLM. Our pipeline separates the generation of candidate reward functions in the generator from the selection of one function in the adjudicator. For the generator, we use LLM-powered evolutionary search to generate a pool of reward functions [7]. In the transparent and customizable adjudicator, we take a new social choice perspective to address the multi-objective nature of our problem: We create a scorer component that evaluates the quality of generated reward functions according to the different objectives (e.g., different prioritization requests). Subsequently, a social welfare function aggregates these “alignment scores” to select the best reward function. The user can select the social welfare function and thereby has additional control over the preferred trade-off between objectives, e.g., maximizing the summed vs. minimum alignment of all objectives. We show that SCLM returns high-quality reward functions even if the computed alignment scores are noisy. In our experiments, we demonstrate that SCLM leads to the selection of reward functions significantly better aligned with complex, multi-objective prompts. Moreover, we also show how it can be used to effectively mitigate the risks of using rewards designed from human prompts: unintended effects for other agents and the ineffective allocation of resources. Overall, SCLM combines the generative power of LLMs to design reward functions with the capabilities of social choice to handle multi-objective decision-making scenarios.
@@ 2 Related Works
The work of Behari et al. [7] is the first to present a Decision-Language Model for generating reward functions for RMABs from human prompts. The model performs a form of evolutionary search to find reward functions aligned with the given prompt in two interleaving phases: generation and reflection. In the generation phase, an LLM generates a set of reward functions. Based on reward function’s performances, in the reflection phase [16, 29], the LLM selects the function best aligned with the prompt. This function is then included in the prompt for the next generation phase or returned. In contrast to our work, DLM mixes generation with selection and does not explicitly account for the multi-objective nature of the reward design problem. Furthermore, in contrast to our work, they focus on small RMAB instances ($\sim 20$ arms). Throughout the paper, we will use a slightly modified variant of DLM adjusted to our setting (see Section C.2) as a baseline.
@@
Research on MORL focuses on learning policies that maximize (and balance between) multiple objective functions, typically via scalarizing the objectives into a single reward function [20] or approximating the Pareto front [28, 31]. In the context of multiagent systems, MORL has been used as a method to ensure the fair treatment of the individual agents [14, 39, 38]. Closest to ours from this line of work is the paper by Zimeng et al. [38], who use ideas from the resource allocation literature to combine multiple objectives into a singular non-linear objective function. However, in contrast to our paper, their focus lies on policy learning for such non-linear objective functions. They neither consider reward design, human preference prompts, LLMs, nor RMABs.
@@
We refer to Appendix A for additional related work.
@@ 3 Preliminaries
An instance of Restless Multi-Armed Bandits (RMAB) is defined by a set of $N$ arms, a time horizon $T$, and a budget $K$. We also refer to arms as agents. Each arm $i\in[N]$ is an independently evolving MDP with state space $\mathcal{S}_{i}$, actions $\mathcal{A}_{i}=\{0,1\}$, transition function $P_{i}:\mathcal{S}_{i}\times\mathcal{A}_{i}\times\mathcal{S}_{i}\to\mathbb{R}_{\geq 0}$, and reward function $R_{i}:\mathcal{S}_{i}\to\mathbb{R}$. We refer to $1$ as the active action corresponding to pulling the arm (i.e., allocating a resource) and 00 as the passive action corresponding to not pulling the arm. We focus on the popular case where each MDP consists of two states, i.e., $\mathcal{S}_{i}=\{0,1\}$ for all $i\in[N]$, yet our methodology applies to MDPs with arbitrary state spaces. We refer to 00 as the bad and $1$ as the good state. For each step in which an agent is in the good state, they derive a utility of $1$, while they derive a utility of 00 in the bad state. Accordingly, agents’ default reward function $R^{*}$ is $R^{*}(s)=s$. We assume that there is a set of categorical features. Each arm is associated with a value of each feature. A global reward function is a reward function defined over features, which induces a reward function for each arm by plugging in its feature values (see Example 1).
@@
For a global reward function $R$, we write $\Pi(R)$ to denote the Whittle index policy for $R$, i.e., the Whittle index policy for the instance where each agent uses the function $R$ after plugging in their feature values as their reward. We refer to $\Pi(R^{*})$ as the default policy. To assess the quality of a global reward function $R$, we often consider the utility feature distribution for some feature $X$. This distribution shows for each value of the feature, the expected utility generated by arms with this feature value under the policy $\Pi(R)$ (see Figure 5(a) for an example).
@@ 4 Problem Statement & Challenges
[…]
@@
To illustrate the tradeoff decisions we face between different clauses when selecting a reward function, in Figure 1, we show two instances from our experiments for a prompt consisting of two prioritization clauses. Every point represents a LLM-designed reward function. The $x$ and $y$ axes represent the quality of the reward function from the perspective of the two prioritized subgroups where higher percentage values indicate more benefits (see Section 6.3 for details). Reward functions marked with stars lie on the Pareto fronts (no other available function dominates them).
In our experiments, we observe that the DLM model from previous work picks functions from very different parts of the Pareto frontier, potentially clearly prioritizing one subgroup over another (see Figure 1(a) for an example). In many other instances, it also picks suboptimal functions, i.e., functions that do not lie on the frontier, that may even harm one of the subgroups while strongly benefiting the other (see Figure 1(b)). This highlights the risks (and shortcomings of DLM) in not accounting for the multi-objective nature of the problem, as it picks reward functions that are inefficient (i.e., dominated) and unfair (i.e., heavily favoring one clause over the other).
@@
Moving from the default reward function to a reward function aligned with a given (prioritization) prompt causes shifts within the distribution of resources and utility. Due to correlations between features, this change might lead to unintended utility shifts for features not specified in the prompt. Figure 5 shows an example of this from our experiments. We present the utility feature distribution for the two features income and education for two reward functions: The reward function selected by DLM for the prompt “Prioritize agents with low income” (orange) and the default reward (blue). While the utility generated by low-income agents increases when moving from the default to the customized reward function, the utility generated by highly educated agents decreases, a side-effect the end-user might be unaware of and that might conflict with their allocation goals. In our proposed approach, we are able to account for this issue by incorporating the prevention of unintended utility shifts as a tradeoff dimension.
@@ 5 Social Choice Language Model (SCLM)
We propose a Social Choice Language Model to generate rewards from human language composite preference prompts (see Figure 3 for a visualization). Separating the generation and selection of reward functions, the model consists of two sequential components. The LLM-powered generator generates a set of candidate reward functions. Subsequently, taking a social-choice-inspired viewpoint, the adjudicator selects a reward function from the pool to be returned to the user in two steps: First, a scorer model computes an alignment score for each reward function with each prioritization clause (i.e., we judge each reward function from the perspective of all relevant “objectives”). Second, a user-defined social welfare function aggregates these scores into a “winning” candidate reward function. By selecting the social welfare function, the user can control the area of the Pareto frontier from which reward functions get selected. While we remark that our model can also be used to tackle multi-objective issues arising when designing rewards in single-agent RL, the details of our components (e.g., the reflection in the generator and the computation of alignment scores) are specific to the multiagent nature of the RMAB problem.
@@ 5.1 Generator
Given a prompt, our generator creates a set of candidate reward functions (as Python code) via a variant of evolutionary search following Behari et al. [7]: We proceed in multiple steps. First, inputting the problem description, feature descriptions and the preference prompt, we ask an LLM to generate code for a reward function. We repeat this query $n_{p}$ times to obtain a set $\mathcal{R}$ of $n_{p}$ candidate reward functions. Afterwards, for each function $R\in\mathcal{R}$ we compute the utility feature distributions of the policy $\Pi(R)$ induced by the reward function $R$ on the given RMAB instance (via repeatedly simulating the policy on the instance). Then, the prompt and the set of candidate reward functions together with the associated utility feature distributions are passed to an LLM, which is asked to select the reward function $R^{\prime}$ from $\mathcal{R}$ best aligned with the prompt [16, 29]. Now, we repeat the whole process, this time including the selected policy $R^{\prime}$ as a seed in the reward function generation prompts. Once we have executed the process $n_{r}$ times, we add all generated $n_{p}\cdot n_{r}$ candidate reward functions $R$ to the pool $\mathcal{R}$ (see Section C.2 for details).
@@ 5.2 Adjudicator
In the following, in Section 5.2.1, we describe how the adjudicator selects the reward function given the scores; in Section 5.2.2, we describe how scores get computed; and in Section 5.2.3, we present a guarantee on the quality of the selected reward function if computed scores are noisy.
@@ 5.2.2 Computing Alignment Scores
For each preference clause $p_{i}\in P$, we compute a reward function $R_{i}$ aligned with $p_{i}$ by casting it as a singular prompt to the DLM pipeline (see Section C.2). For each $R\in\mathcal{R}$, we compute as $s_{i}(R)$ the expected reward according to reward function $R_{i}$ produced by policy $\Pi(R)$ (again, we approximate this quantity by running multiple simulations). Accordingly, $s_{i}(R)$ quantifies the quality of the policy induced by the candidate reward function $R$ from the perspective of $p_{i}$ (as captured by $R_{i}$). As the scale of the reward functions can vary significantly among preference clauses, we normalize the scores by the performance of the default policy, i.e., we compute $\frac{s_{i}(R)-s_{i}(R^{*})}{s_{i}(R^{*})}$.
@@
The Simulator Scorer Model assumes access to reward functions capturing individual preference clauses well. If no well-aligned reward functions can be obtained, the performance of SCLM-SIM can deteriorate because it can become noisy. Another disadvantage of SCLM-SIM is that the scores in SCLM-SIM are all computed via simulation, which can become computationally costly. Motivated by this, we propose a quicker and more flexible LLM-based approach, where we prompt an LLM to rate the alignment of a candidate reward function with a preference clause. In particular, for each $R\in\mathcal{R}$ and $p_{i}\in P$, we use a prompt that includes $R$, $p_{i}$, and the utility feature distributions produced by policy $\Pi(R)$. We ask the LLM to rank how well $R$ aligns with the preference clause $p_{i}$ on a scale from 1 to 5 (see Appendix F for prompt texts).
@@
Aligning reward functions to a prioritization prompt may cause (unintended) utility shifts in other features (e.g., due to feature correlations, shifting utility to low-income beneficiaries might shift it away from more educated ones). See Figure 5 for a concrete example. SCLM offers users the option to explicitly prevent these shifts by adding additional clauses (“objectives”) to the prompt: Given a prompt $P$ (e.g., the prompt from Example 1), for each feature not referenced in the prompt, the user can add a new preference clause requesting a minimum shift in the utility distribution of this feature (e.g., for Example 1 they could add “do not change the utility distribution for feature $C$”). To compute the alignment score $s_{i}(R)$ between a reward function $R$ and a clause $p_{i}$=“minimize utility shift for feature $X$”, we compare feature $X$’s utility distribution under the default policy with its utility distribution under the policy $\Pi(R)$. Specifically, we quantify the difference using the Earth mover’s distance (EMD) between the two distributions. Afterward, we apply 00-$1$ normalization to all scores $s_{i}(R)_{R\in\mathcal{R}}$ for prompt $p_{i}$, which are input to the social welfare function (along with the alignment scores for the other clauses).
Another potential risk of aligning a reward function with a prioritization prompt is that it can sharply decrease the summed utility generated by all agents: The user might request the prioritization of a subpopulation that does not benefit much from receiving a resource, leading to severe drops in the summed utility generated by all agents. Users can address this issue in our model by adding a clause $p_{i}$=“maximize the total generated utility” to the prompt. As the alignment score $s_{i}(R)$ of $p_{i}$ with some reward function $R$ we compute the summed utility, i.e., the total number of steps in which arms are in an active state, generated by all agents under the policy $\Pi(R)$ (computed via multiple simulations of the policy on the given instance). We again apply 00-$1$ normalization to all scores $s_{i}(R)_{R\in\mathcal{R}}$ for prompt $p_{i}$.

### v5

@@ 1 Introduction
Unfortunately, handcrafting reward functions is often a challenging and time-consuming task for humans because of the complex relationship between reward functions and policy outcomes [12, 8, 17]. Further, the multiagent nature of the RMAB problem adds a new twist to the problem of reward design in RL: It becomes fundamentally multi-objective. Consider the above example prompt asking for the prioritization of two subpopulations. As these subpopulations may contain different agents, selecting a reward function will most likely involve trading off the interests of the low-income vs. older beneficiaries, making this a multi-objective problem. If this multi-objective nature is ignored, a selected reward function might heavily favor one of the two groups (e.g., leading to the allocation of many resources to low-income beneficiaries, and no resources to older ones).
This problem of multi-objective reward function modification even extends beyond Restless Multi-Armed Bandits, to challenges relevant in using Game Theory and AI for security, for instance, in Stackelberg Security Games (SSGs) applications [33]. In SSGs, optimal security strategies are highly sensitive to the defined payoff functions, which typically reflect the defender’s and attacker’s objectives. Using a natural language interface offers defenders to refine and balance these payoff functions when faced with conflicting objectives can act as a powerful tool. This is in contrast with traditional iterative and manual adjustments presented in prior SSG applications. For instance, the ARMOR system deployed at LAX airport for security resource allocation [26] originally included a graphical user interface to enable payoff adjustments, but this interface was rudimentary. A natural language interface, as proposed in our work, could significantly enhance the practical utility of SSGs by allowing security personnel to express nuanced objectives (e.g., "prioritize slightly more air marshals on routes to Tokyo and Paris" using a system like IRIS [34]) directly, and have these preferences be translated into balanced and effective security strategies. This capability addresses a critical need for more flexible and human-centric control over complex security resource allocation.
@@
To our knowledge, we are the first to address the multi-objective nature of LLM-powered reward design in RMABs in particular and multiagent planners in general. Closest to our paper is the work by Behari et al. [7] who proposed a fully LLM-based Decision-Language Model for RMABs to generate and select reward functions (as code) from human language prompts. However, as argued in Sections˜2 and 4, the DLM model is not properly equipped to handle the multi-objective nature of the problem, as the LLM selects functions in an unpredictable, hard-to-control and sometimes (clearly) suboptimal way that does not adequately take into account and balance the different objectives.
@@
We present a Social Choice Language Model (SCLM) that designs reward functions (as Python code) aligned with complex, multi-objective human language preferences; see Figure˜3 for an overview of SCLM. Our pipeline separates the generation of candidate reward functions in the generator from the selection of one function in the adjudicator. For the generator, we use LLM-powered evolutionary search to generate a pool of reward functions [7]. In the transparent and customizable adjudicator, we take a new social choice perspective to address the multi-objective nature of our problem: We create a scorer component that evaluates the quality of generated reward functions according to the different objectives (e.g., different prioritization requests). Subsequently, a social welfare function aggregates these “alignment scores” to select the best reward function. The user can select the social welfare function and thereby has additional control over the preferred trade-off between objectives, e.g., maximizing the summed vs. minimum alignment of all objectives. We show that SCLM returns high-quality reward functions even if the computed alignment scores are noisy. In our experiments, we demonstrate that SCLM leads to the selection of reward functions significantly better aligned with complex, multi-objective prompts. Moreover, we also show how it can be used to effectively mitigate the risks of using rewards designed from human prompts: unintended effects for other agents and the ineffective allocation of resources. Overall, SCLM combines the generative power of LLMs to design reward functions with the capabilities of social choice to handle multi-objective decision-making scenarios.
@@ 2 Related Works
The work of Behari et al. [7] is the first to present a Decision-Language Model for generating reward functions for RMABs from human prompts. The model performs a form of evolutionary search to find reward functions aligned with the given prompt in two interleaving phases: generation and reflection. In the generation phase, an LLM generates a set of reward functions. Based on reward function’s performances, in the reflection phase [17, 31], the LLM selects the function best aligned with the prompt. This function is then included in the prompt for the next generation phase or returned. In contrast to our work, DLM mixes generation with selection and does not explicitly account for the multi-objective nature of the reward design problem. Furthermore, in contrast to our work, they focus on small RMAB instances ($\sim 20$ arms). Throughout the paper, we will use a slightly modified variant of DLM adjusted to our setting (see Section˜C.2) as a baseline.
@@
Research on MORL focuses on learning policies that maximize (and balance between) multiple objective functions, typically via scalarizing the objectives into a single reward function [21] or approximating the Pareto front [30, 35, 15]. In the context of multiagent systems, MORL has been used as a method to ensure the fair treatment of the individual agents [14, 43, 42]. Closest to ours from these lines of work are the papers by [42] and [15]. [42] uses ideas from the resource allocation literature to combine multiple objectives into a singular non-linear objective function and focuses on policy learning for such non-linear objective functions. [15] focuses on finding a set of policies that approximate the Pareto front for various sequential planning problems. However, in contrast to our paper, neither considers reward design, human natural language preference prompts and LLMs.
@@
We refer to Appendix˜A for additional related work.
@@ 3 Preliminaries
An instance of Restless Multi-Armed Bandits (RMAB) is defined by a set of $N$ arms, a time horizon $T$, and a budget $K$. We also refer to arms as agents. Each arm $i\in[N]$ is an independently evolving MDP with state space $\mathcal{S}_{i}$, actions $\mathcal{A}_{i}=\{0,1\}$, transition function $P_{i}:\mathcal{S}_{i}\times\mathcal{A}_{i}\times\mathcal{S}_{i}\to\mathbb{R}_{\geq 0}$, and reward function $R_{i}:\mathcal{S}_{i}\to\mathbb{R}$. We refer to $1$ as the active action corresponding to pulling the arm (i.e., allocating a resource) and 0 as the passive action corresponding to not pulling the arm. We focus on the popular case where each MDP consists of two states, i.e., $\mathcal{S}_{i}=\{0,1\}$ for all $i\in[N]$, yet our methodology applies to MDPs with arbitrary state spaces. We refer to 0 as the bad and $1$ as the good state. For each step in which an agent is in the good state, they derive a utility of $1$, while they derive a utility of 0 in the bad state. Accordingly, agents’ default reward function $R^{*}$ is $R^{*}(s)=s$. We assume that there is a set of categorical features. Each arm is associated with a value of each feature. A global reward function is a reward function defined over features, which induces a reward function for each arm by plugging in its feature values (see Example˜1).
@@
For a global reward function $R$, we write $\Pi(R)$ to denote the Whittle index policy for $R$, i.e., the Whittle index policy for the instance where each agent uses the function $R$ after plugging in their feature values as their reward. We refer to $\Pi(R^{*})$ as the default policy. To assess the quality of a global reward function $R$, we often consider the utility feature distribution for some feature $X$. This distribution shows for each value of the feature, the expected utility generated by arms with this feature value under the policy $\Pi(R)$ (see Figure˜5(a) for an example).
@@ 4 Problem Statement & Challenges
Consider an RMAB instance with three binary features $A$, $B$, and $C$. A preference prompt $P$ could be “Prioritize agents with $A=0$ and prioritize agents with B=1B=1”, i.e., $P=\{$“prioritize agents with $A=0$, “prioritize agents with B=1B=1}\}. Two possible global reward functions for the prompt are $R^{\prime}(s)=s\cdot(1-A)\cdot B$ and $R^{\prime\prime}(s)=s\cdot(1-A)+s\cdot B$. For function $R^{\prime\prime}$, the reward of an agent $i$ with $A=0$ and $B=1$ is $R_{i}(s)=2s$, while the reward of an agent $j$ with $A=1$ and $B=1$ is $R_{j}(s)=s$. Selecting $R^{\prime\prime}$, agent $i$ is more likely to receive a resource than agent $j$, as the good state contributes more reward for $i$.
@@
To illustrate the tradeoff decisions we face between different clauses when selecting a reward function, in Figure˜1, we show two instances from our experiments for a prompt consisting of two prioritization clauses. Every point represents a LLM-designed reward function. The $x$ and $y$ axes represent the quality of the reward function from the perspective of the two prioritized subgroups where higher percentage values indicate more benefits (see Section˜6.3 for details). Reward functions marked with stars lie on the Pareto fronts (no other available function dominates them).
In our experiments, we observe that the DLM model from previous work picks functions from very different parts of the Pareto frontier, potentially clearly prioritizing one subgroup over another (see Figure˜1(a) for an example). In many other instances, it also picks suboptimal functions, i.e., functions that do not lie on the frontier, that may even harm one of the subgroups while strongly benefiting the other (see Figure˜1(b)). This highlights the risks (and shortcomings of DLM) in not accounting for the multi-objective nature of the problem, as it picks reward functions that are inefficient (i.e., dominated) and unfair (i.e., heavily favoring one clause over the other).
@@
Moving from the default reward function to a reward function aligned with a given (prioritization) prompt causes shifts within the distribution of resources and utility. Due to correlations between features, this change might lead to unintended utility shifts for features not specified in the prompt. Figure˜5 shows an example of this from our experiments. We present the utility feature distribution for the two features income and education for two reward functions: The reward function selected by DLM for the prompt “Prioritize agents with low income” (orange) and the default reward (blue). While the utility generated by low-income agents increases when moving from the default to the customized reward function, the utility generated by highly educated agents decreases, a side-effect the end-user might be unaware of and that might conflict with their allocation goals. In our proposed approach, we are able to account for this issue by incorporating the prevention of unintended utility shifts as a tradeoff dimension.
@@ 5 Social Choice Language Model (SCLM)
We propose a Social Choice Language Model to generate rewards from human language composite preference prompts (see Figure˜3 for a visualization). Separating the generation and selection of reward functions, the model consists of two sequential components. The LLM-powered generator generates a set of candidate reward functions. Subsequently, taking a social-choice-inspired viewpoint, the adjudicator selects a reward function from the pool to be returned to the user in two steps: First, a scorer model computes an alignment score for each reward function with each prioritization clause (i.e., we judge each reward function from the perspective of all relevant “objectives”). Second, a user-defined social welfare function aggregates these scores into a “winning” candidate reward function. By selecting the social welfare function, the user can control the area of the Pareto frontier from which reward functions get selected. While we remark that our model can also be used to tackle multi-objective issues arising when designing rewards in single-agent RL, the details of our components (e.g., the reflection in the generator and the computation of alignment scores) are specific to the multiagent nature of the RMAB problem.
@@ 5.1 Generator
Given a prompt, our generator creates a set of candidate reward functions (as Python code) via a variant of evolutionary search following Behari et al. [7]: We proceed in multiple steps. First, inputting the problem description, feature descriptions and the preference prompt, we ask an LLM to generate code for a reward function. We repeat this query $n_{p}$ times to obtain a set $\mathcal{R}$ of $n_{p}$ candidate reward functions. Afterwards, for each function $R\in\mathcal{R}$ we compute the utility feature distributions of the policy $\Pi(R)$ induced by the reward function $R$ on the given RMAB instance (via repeatedly simulating the policy on the instance). Then, the prompt and the set of candidate reward functions together with the associated utility feature distributions are passed to an LLM, which is asked to select the reward function $R^{\prime}$ from $\mathcal{R}$ best aligned with the prompt [17, 31]. Now, we repeat the whole process, this time including the selected policy $R^{\prime}$ as a seed in the reward function generation prompts. Once we have executed the process $n_{r}$ times, we add all generated $n_{p}\cdot n_{r}$ candidate reward functions $R$ to the pool $\mathcal{R}$ (see Section˜C.2 for details).
@@ 5.2 Adjudicator
In the following, in Section˜5.2.1, we describe how the adjudicator selects the reward function given the scores; in Section˜5.2.2, we describe how scores get computed; and in Section˜5.2.3, we present a guarantee on the quality of the selected reward function if computed scores are noisy.
@@ 5.2.2 Computing Alignment Scores
For each preference clause $p_{i}\in P$, we compute a reward function $R_{i}$ aligned with $p_{i}$ by casting it as a singular prompt to the DLM pipeline (see Section˜C.2). For each $R\in\mathcal{R}$, we compute as $s_{i}(R)$ the expected reward according to reward function $R_{i}$ produced by policy $\Pi(R)$ (again, we approximate this quantity by running multiple simulations). Accordingly, $s_{i}(R)$ quantifies the quality of the policy induced by the candidate reward function $R$ from the perspective of $p_{i}$ (as captured by $R_{i}$). As the scale of the reward functions can vary significantly among preference clauses, we normalize the scores by the performance of the default policy, i.e., we compute $\frac{s_{i}(R)-s_{i}(R^{*})}{s_{i}(R^{*})}$.
@@
The Simulator Scorer Model assumes access to reward functions capturing individual preference clauses well. If no well-aligned reward functions can be obtained, the performance of SCLM-SIM can deteriorate because it can become noisy. Another disadvantage of SCLM-SIM is that the scores in SCLM-SIM are all computed via simulation, which can become computationally costly. Motivated by this, we propose a quicker and more flexible LLM-based approach, where we prompt an LLM to rate the alignment of a candidate reward function with a preference clause. In particular, for each $R\in\mathcal{R}$ and $p_{i}\in P$, we use a prompt that includes $R$, $p_{i}$, and the utility feature distributions produced by policy $\Pi(R)$. We ask the LLM to rank how well $R$ aligns with the preference clause $p_{i}$ on a scale from 1 to 5 (see Appendix˜F for prompt texts).
@@
Aligning reward functions to a prioritization prompt may cause (unintended) utility shifts in other features (e.g., due to feature correlations, shifting utility to low-income beneficiaries might shift it away from more educated ones). See Figure˜5 for a concrete example. SCLM offers users the option to explicitly prevent these shifts by adding additional clauses (“objectives”) to the prompt: Given a prompt $P$ (e.g., the prompt from Example˜1), for each feature not referenced in the prompt, the user can add a new preference clause requesting a minimum shift in the utility distribution of this feature (e.g., for Example˜1 they could add “do not change the utility distribution for feature CC”). To compute the alignment score $s_{i}(R)$ between a reward function $R$ and a clause $p_{i}$=“minimize utility shift for feature XX”, we compare feature $X$’s utility distribution under the default policy with its utility distribution under the policy $\Pi(R)$. Specifically, we quantify the difference using the Earth mover’s distance (EMD) between the two distributions. Afterward, we apply 0-11 normalization to all scores $s_{i}(R)_{R\in\mathcal{R}}$ for prompt $p_{i}$, which are input to the social welfare function (along with the alignment scores for the other clauses).
Another potential risk of aligning a reward function with a prioritization prompt is that it can sharply decrease the summed utility generated by all agents: The user might request the prioritization of a subpopulation that does not benefit much from receiving a resource, leading to severe drops in the summed utility generated by all agents. Users can address this issue in our model by adding a clause $p_{i}$=“maximize the total generated utility” to the prompt. As the alignment score $s_{i}(R)$ of $p_{i}$ with some reward function $R$ we compute the summed utility, i.e., the total number of steps in which arms are in an active state, generated by all agents under the policy $\Pi(R)$ (computed via multiple simulations of the policy on the given instance). We again apply 0-11 normalization to all scores $s_{i}(R)_{R\in\mathcal{R}}$ for prompt $p_{i}$.